Ticket Number: HEA-NOC/20050413-8 Ticket Status: CLOSE
Ticket Type: scheduled Resolver: HEA-NOC
Ticket Opened: 20050413 12:27 UTC+1 Problem Start: 20050413 13:30 UTC+1
Ticket Closed: 20050418 09:53 UTC+1 Problem End: 20050413 14:00 UTC+1
IRCSET will be upgrading their connection to the network. This will entail
a change of bearer from serial line to ethernet, and will result in a
small amount of downtime.
Traffic to and from IRCSET will be interrupted
20050413 12:17 BB Scheduling work
20050418 09:52 BB Recabling completed, new router installed. Kerna have
reconfigured IRCSET switching and firewalling to use the new connection.
Time to Fix:
This ticket can be monitored at http://www.hea.net/tickets/20050413-8